1. A tank contains helium for filling balloons. It holds 6.75 L of He at 42.2 atm and 22.9 °C. What would be the volume of this helium at 2.39 atm and 13.2 °C?

2. The volume of a sample of ethane, C2H6, is 2.38 L at 452 torr and 30 °C. What volume will it occupy at standard temperature and pressure (STP)?
1
Expert's answer
2019-10-26T02:58:37-0400

PV=nRT

1 atm = 101325 Pa

0 C = 273.15K


1) P1V1/P2V2=T1/T2

42.2*6.75/2.39/V2=(22.9+271.15)/(13.2+273.15)

V2=115.3 L

2) STP sunce 1982 T=273.15K P = 100000Pa

452 torr = 60261.54

60261.54* 2.38/ 100000/V2= (30+273.15)/273.15

V2=1.29 L